In a student's French class, the final grade is based on six categories.

The categories, grades, and weights are shown in this table.

CategoryGradeWeight %

Test 15216

Test25820

Test 37018

Homework716

Semester Project907

Final Exam7133

Compute a weighted average to determine the student's overall final grade in the course. Record the overall final grade below as a percentage. Round accurately to two decimal places.

Final Grade: %
Answer #1

Answer:

Given,

In a student's French class, the final grade is based on six categories

To give the overall final grade

Final grade = xi*P(xi)

Final grade = 152*0.16 + 258*0.20 + 370*0.18 + 71*0.6 + 90*0.7 + 71*0.33

= 24.32 + 51.6 + 66.6 + 42.6 + 63 + 23.43

Final grade = 271.55